please help me with the following linear equation with decimals:
x-0.08x=460
and 2x/2 = x

Answer by mananth(16949)

x-0.08x=460
1.00-0.08 = 0.92
0.92x=460
x=460/0.92
x= 460 * 100 / 0.92 * 100
x= 46000 / 92
x=500